A light-bulb manufacturer regularly advertises that his bulbs last 900 hours with a standard deviation of 75 hours. A random sample is chosen before each campaign to make sure that the claim is correct. A sample of 200 bulbs shows a mean of 925 hours. Can the advertising’s claim be considered an underestimate at the 0.05 level of significance?
1
Expert's answer
2020-04-01T15:28:23-0400

"Ho: \mu=900\\
Ha: \mu>900"

Test statistic is given by

"\frac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}="

"=\frac{925-900}{\frac{75}{\sqrt{200}}}=4.714"

From z table or =NORM.S.INV(0.95) Excel formula, the critical value is 1.645. Since test statistic is greater than the critical value, we reject the null hypothesis and conclude that the advertising claim is an underestimate.